two questions available please answer both

answer both!!!

Proctuc thco piant will reject a boten Syringes if two or more cdefece singes srinojes taxen Frcm tha batch. SUppcse the baton contains S ctefecive stringes of Make histegrom sncuina the Prcbabites of r=0山2.3 , il and 12 cleFeenve Syringes in a randcon sam pa oc เ2Syring :una+-is +hL expected # of clefeen synnges your inspec tor will ind 2 be accep tecl ? d ind O
5). Does the kidtacei, mace a devenaa ?邛You ove aiking Yes abcot pho toogropnthe answer Ages of chuccn in Hcusenowl under 7- of uS nousnads that buy um o 30 % Part uncder the aoe of a s Let vepresent the t of such nouse naids +ho buy m shcwing the ind the mean and stcndord cdeicunam house bsio bi suppose that housenciols are chosem te no chuaren uncler al yeaxs clal Let r represent the number cf Sucn hsenelals thot bu him he Find the mean 3 stendord devion om
Answer #1

Ans:

a)Use binomial distribution with n=12 and p=0.05

P(x=r)=12Cr*0.05r*0.9512-r

r P(r)
0 0.5404
1 0.3413
2 0.0988
3 0.0173
4 0.0021
5 0.0002
6 0.0000
7 0.0000
8 0.0000
9 0.0000
10 0.0000
11 0.0000
12 0.0000

P(r) 0.6000 0.5000 0.3000 0.2000 0.1000 0.0000 1 23 4 56 78 9 10 11 12 13

b)

Expected umber of defective syringes,mu=np=12*0.05=0.6

c)P(rejected)=P(x>=2)

P(accepted)=P(x<=1)

=P(x=0)+P(x=1)

=0.95^12+12C1*0.05*0.9511

=0.5403+0.3413

=0.8816

d)

standard deviation,sigma=sqrt(npq)=sqrt(12*0.05*0.95)=0.755